OpenStudy (anonymous):

Choose 2 axioms that allows 22 + (m + 8) to be written as m + 30

OpenStudy (anonymous):

Communitivity and Associativity I believe ( i know i spelled those wrong) Because addition is communitive: \[22+(m+8)=(m+8)+22\]

OpenStudy (anonymous):

Then because addition is associative: \[(m+8)+22=m+(8+22) = m+30\]

OpenStudy (anonymous):

a+(b+c) = (a+b)+c a+b+c = a+(b+c)